# Rationality based Innate-Values-driven Reinforcement Learning

### v1

@@ Abstract
Innate values describe agents’ intrinsic motivations, which reflect their inherent interests and preferences to pursue goals and drive them to develop diverse skills satisfying their various needs. The essence of reinforcement learning (RL) is learning from interaction based on reward-driven behaviors, much like natural agents. It is an excellent model to describe the innate-values-driven (IV) behaviors of AI agents. Especially developing the awareness of the AI agent through balancing internal and external utilities based on its needs in different tasks is a crucial problem for individuals learning to support AI agents integrating human society with safety and harmony in the long term. This paper proposes a hierarchical compound intrinsic value reinforcement learning model – innate-values-driven reinforcement learning termed IVRL to describe the complex behaviors of AI agents’ interaction. We formulated the IVRL model and proposed two IVRL models: DQN and A2C. By comparing them with benchmark algorithms such as DQN, DDQN, A2C, and PPO in the Role-Playing Game (RPG) reinforcement learning test platform VIZDoom, we demonstrated that rationally organizing various individual needs can effectively achieve better performance.
1 Introduction
@@
In natural systems, motivation is concerned explicitly with the activities of creatures that reflect the pursuit of a particular goal and form a meaningful unit of behavior in this function heckhausen2018motivation . Furthermore, intrinsic motivations describe incentives relating to an activity itself, and these incentives residing in pursuing an activity are intrinsic. Intrinsic motivations deriving from an activity may be driven primarily by interest or activity-specific incentives, depending on whether the object of an activity or its performance provides the main incentive schiefele1996motivation . They also fall in the category of cognitive motivation theories, which include theories of the mind that tend to be abstracted from the biological system of the behaving organism merrick2013novelty .
However, when we analyze natural agents, such as humans, they are usually combined motivation entities. They have biological motivations, including physiological, safety, and existence needs; social motivation, such as love and esteem needs; and cognitive motivation, like self-actualization or relatedness and growth needs merrick2009motivated . The combined motivation theories include Maslow’s Hierarchy of Needs maslow1958dynamic and Alderfer’s Existence Relatedness Growth (ERG) theory alderfer1972existence . Fig. 2 and 2 illustrate the general innate values (intrinsic motivations) model and various models with three-level needs of different amounts, respectively.
@@
Many researchers regard motivated behavior as behavior that involves the assessment of the consequences of behavior through learned expectations, which makes motivation theories tend to be intimately linked to theories of learning and decision-making baldassarre2013intrinsically . In particular, intrinsic motivation leads organisms to engage in exploration, play, strategies, and skills driven by expected rewards. The computational theory of reinforcement learning (RL) addresses how predictive values can be learned and used to direct behavior, making RL naturally relevant to studying motivation.
Figure 1: The illustration of the proposed innate-values-driven reinforcement learning (IVRL) model.
Figure 2: The illustration of innate values models with three-level needs of different amounts. S: Small; M: Medium; L: Large
@@
In artificial intelligence, researchers propose various abstract computational structures to form the fundamental units of cognition and motivations, such as states, goals, actions, and strategies. For intrinsic motivation modeling, the approaches can be generally classified into three categories: prediction-based schmidhuber1991curious ; schmidhuber2010formal , novelty-based marsland2000real ; merrick2009motivated , and competence-based barto2004intrinsically ; schembri2007evolution . Furthermore, the concept of intrinsic motivation was introduced in machine learning and robotics to develop artificial systems learning diverse skills autonomously yang2024bayesian . The idea is that intelligent machines and robots could autonomously acquire skills and knowledge under the guidance of intrinsic motivations and later exploit such knowledge and skills to accomplish tasks more efficiently and faster than if they had to acquire them from scratch baldassarre2013intrinsically .
@@
In other words, by investigating intrinsically motivated learning systems, we would clearly improve the utility and autonomy of intelligent artificial systems in dynamic, complex, and dangerous environments yang2022game ; yang2023hierarchical . Specifically, compared with the traditional RL model, intrinsically motivated RL refines it by dividing the environment into an external environment and an internal environment, which clearly generates all reward signals within the organism [^1] Here, the organism represents all the components of the internal environment in the AI agent. baldassarre2013intrinsically . However, although the extrinsic reward signals are triggered by the objects and events of the external environment, and activities of the internal environment cause the intrinsic reward signals, it is hard to determine the complexity and variability of the intrinsic rewards (innate values) generating mechanism.
@@
To address those gaps, we introduce the innate-values-driven reinforcement learning (IVRL) model to describe the complex behaviors in AI agents’ interactions by integrating with combined motivation theories. We formalize the idea and propose two IVRL models based on classic DQN and A2C algorithms. Then, we compare them with benchmark RL algorithms such as DQN mnih2015human , DDQN wang2016dueling , A2C mnih2016asynchronous , and PPO schulman2017proximal in the RPG RL test platform VIZDoom Kempka2016ViZDoom ; Wydmuch2019ViZdoom . The results demonstrate that the IVRL model can achieve convergence and adapt efficiently to complex and challenging tasks.
2 Approach Overview
@@
We assume that all the AI agents (like robots) interact in the same working scenario, and their external environment includes all the other group members and mission setting. In contrast, the internal environment consists of individual perception components including various sensors (such as Lidar and camera), the critic module involving intrinsic motivation analysis and innate values generation, the RL brain making the decision based on the feedback of rewards and description of the current state (including internal and external) from the critic module, and actuators relating to all the manipulators and operators executing the RL brain’s decisions as action sequence and strategies. Fig. 4 illustrates the proposed IVRL model.
Compared with the traditional RL model, our model generates the input state and rewards from the critic module instead of directly from the environment. This means that the individual needs to calculate rewards based on the innate value and current utilities and then update its current internal status – the hierarchical needs model. For example, supposing two agents $a^{1}$ and $a^{2}$ have different innate value models: Fig. 2 S-M-L and L-M-S. We use health points, energy levels, and task achievement to represent their safety needs $n_{s}$ (low-level intrinsic motivations), basic needs $n_{b}$ (middle-level intrinsic motivations), and teaming needs $n_{t}$ (High-level intrinsic motivations) yang2020hierarchical ; yang2020needs ; yang2021can , respectively. Considering $n_{s}^{1}>n_{s}^{2}$, $n_{b}^{1}=n_{b}^{2}$, and $n_{t}^{1}<n_{t}^{2}$, if $a^{1}$ and $a^{2}$ receive the external repairing signal simultaneously, the innate rewards $r^{1}$ will larger than $r^{2}$ based on their innate value model (low-level safety needs $n_{s}^{1}>n_{s}^{2}$). In contrast, if they get the order to collaborate with other agents fulfilling a task (high-level teaming needs $n_{t}^{1}<n_{t}^{2}$) at the same time, the agent $a^{2}$ will receive more credits from it ($r^{1}<r^{2}$). In other words, due to different innate value models, their intrinsic motivations present significant differences, which will lead to different innate rewards for performing the same task.
@@
Specifically, we formalize the IVRL of an AI agent with an external environment using a Markov decision process (MDP) puterman2014markov . The MDP is defined by the tuple $\langle\mathcal{S,A,R,T,}\gamma\rangle$ where $\mathcal{S}$ represents the finite sets of internal state $S_{i}$ [^2] The internal state $S_{i}$ describes an agent’s innate value distribution and presents the dominant intrinsic motivation based on the external state $S_{e}$. and external states $S_{e}$. $\mathcal{A}$ represents a finite set of actions. The transition function $\mathcal{T}$: $\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow$ [0, 1] determines the probability of a transition from any state $s\in\mathcal{S}$ to any state $s^{\prime}\in\mathcal{S}$ given any possible action $a\in\mathcal{A}$. Assuming the critic function is $\mathcal{C}$, which describes the individual innate value model. The reward function $\mathcal{R}=\mathcal{C}(S_{e}):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ defines the immediate and possibly stochastic innate reward $\mathcal{C}(S_{e})$ that an agent would receive given that the agent executes action $a$ which in state $s$ and it is transitioned to state $s^{\prime},\gamma\in[0,1)$ the discount factor that balances the trade-off between innate immediate and future rewards.
@@
Figure 4: The illustration of the innate-values-driven reinforcement learning (IVRL) model.
2.1 The Source of Randomness
In our model, the randomness comes from three sources. The randomness in action is from the policy function: $A\sim\pi(\cdot|s)$; the needs weight function: $W\sim\omega(\cdot|s)$ makes the randomness of innate values; the state-transition function: $S^{\prime}\sim p(\cdot|s,a)$ causes the randomness in state.
Supposing at current state $s_{t}$ an agent has a needs weight matrix $N_{t}$ (Eq. (1)) in a mission, which presents its innate value weights for different levels of needs. Correspondingly, it has a utility matrix $U_{t}$ (Eq. (1)) for specific needs resulting from action $a_{t}$. Then, we can calculate its reward $R_{t}$ for $a_{t}$ through Eq. (2) at the state $s_{t}$.
$N_{t}=\left[\begin{matrix}n_{11}&n_{12}&\cdots&n_{1m}\\ n_{21}&n_{22}&\cdots&n_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ n_{n1}&n_{n2}&\cdots&n_{nm}\\ \end{matrix}\right];~{}~{}~{}U_{t}=\left[\begin{matrix}u_{11}&u_{12}&\cdots&u_{1m}\\ u_{21}&u_{22}&\cdots&u_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ u_{n1}&u_{n2}&\cdots&u_{nm}\\ \end{matrix}\right]$ (1)
@@ 2.1 The Source of Randomness
In the process, the agent will first generate the needs weight and action based on the current state, then, according to the feedback utilities and the needs weights, calculate the current reward and iterate the process until the end of an episode. Fig. 3 illustrates the trajectory of state $S$, needs weight $W$, action $A$, and reward $R$ in the IVRL model.
2.2 Randomness in Discounted Returns
According to the above discussion, we define the discounted return $G$ at $t$ time as cumulative discounted rewards in the IVRL model (Eq. (3)) and $\gamma$ is the discount factor.
@@ 2.2 Randomness in Discounted Returns
2.3 Action-Innate-Value Function
Based on the discounted return Eq. (3) and its random factors – Eq. (4), (5), and (6), we can define the Action-Innate-Value function as the expectation of the discounted return $G$ at $t$ time (Eq. (7)).
@@ 2.3 Action-Innate-Value Function
2.4 State-Innate-Value Function
Furthermore, we can define the State-Innate-Value function as Eq. (8), which calculates the expectation of $Q_{\pi,\omega}(s_{t},w_{t},a_{t})$ for action $A$ and reflects the situation in the state $s_{t}$ with the innate value judgment $w_{t}$.
@@ 2.4 State-Innate-Value Function
2.5 Approximate the Action-Innate-Value Function
The agent’s goal is to interact with the environment by selecting actions to maximize future rewards based on its innate value judgment. We make the standard assumption that a factor of $\gamma$ per time-step discounts future rewards and define the future discounted return at time $t$ as Eq. (3). Moreover, we can define the optimal action-value function $Q^{*}(s,a,w)$ as the maximum expected return achievable by following any strategy after seeing some sequence $s$, making corresponding innate value judgment $w$, and then taking action $a$, where $\omega$ is a needs weight function describing sequences about innate value weights and $\pi$ is a policy mapping sequences to actions.
@@ 2.5 Approximate the Action-Innate-Value Function
Furthermore, the same as the DQN mnih2015human , we use a function approximator (Eq. (11)) to estimate the action-innate-value function.
@@
Figure 5: Illustration of the action-innate-value network generating Needs-Behavior distribution.
@@
Where Eq. 13 is the target for iteration $i$ and $\sigma(s,w,a)$ is a probability distribution over sequences $s$, needs weights $w$, and action $a$ that we refer to as the needs-behavior distribution. We can approximate the gradient as follows:
$\begin{split}\nabla_{\theta_{i}}L_{i}(\theta_{i})=\mathbb{E}_{s,w,a\sim\sigma(\cdot);s^{\prime}\sim\epsilon}\bigg{[}\bigg{(}r+\gamma\max_{w^{\prime},a^{\prime}}Q(s^{\prime},w^{\prime},a^{\prime};\theta_{i-1})-Q(s,w,a;\theta_{i})\bigg{)}\nabla_{\theta_{i}}Q(s,w,a;\theta_{i})\bigg{]}\end{split}$ (14)
Instead of computing the full expectations in the Eq. (14), stochastic gradient descent is usually computationally expedient to optimize the loss function. Here, the weights $\theta$ are updated after every time step, and single samples from the needs-behavior distribution $\sigma$ and the emulator $\epsilon$ replace the expectations, respectively.

### v2

# Innate-Values-driven Reinforcement Learning Based Cognitive Modeling

@@ Abstract
Innate values describe agents’ intrinsic motivations, which reflect their inherent interests and preferences for pursuing goals and drive them to develop diverse skills that satisfy their various needs. Traditional reinforcement learning (RL) is learning from interaction based on the feedback rewards of the environment. However, in real scenarios, the rewards are generated by agents’ innate value systems, which differ vastly from individuals based on their needs and requirements. In other words, considering the AI agent as a self-organizing system, developing its awareness through balancing internal and external utilities based on its needs in different tasks is a crucial problem for individuals learning to support others and integrate community with safety and harmony in the long term. To address this gap, we propose a new RL model termed innate-values-driven RL (IVRL) based on combined motivations’ models and expected utility theory to mimic its complex behaviors in the evolution through decision-making and learning. Then, we introduce two IVRL-based models: IV-DQN and IV-A2C. By comparing them with benchmark algorithms such as DQN, DDQN, A2C, and PPO in the Role-Playing Game (RPG) reinforcement learning test platform VIZDoom, we demonstrated that the IVRL-based models can help the agent rationally organize various needs, achieve better performance effectively.
I Introduction
@@
In natural systems, motivation is concerned explicitly with the activities of creatures that reflect the pursuit of a particular goal and form a meaningful unit of behavior in this function [1]. From the neuroscience perspective, intrinsic motivation refers to an agent’s spontaneous tendencies to be curious and interested, to seek out challenges, and to exercise and develop their skills and knowledge, even without operationally separable rewards [2]. Furthermore, they describe incentives relating to an activity itself, and these incentives residing in pursuing an activity are intrinsic [3]. Moreover, intrinsic motivations deriving from an activity may be driven primarily by interest or activity-specific incentives, depending on whether the object of an activity or its performance provides the main incentive [4]. They also fall in the category of cognitive motivation theories, which include theories of the mind that tend to be abstracted from the biological system of the behaving organism [5].
Figure 1: The illustration five human agents with various personalities presenting different amounts of innate values and preferences based on five levels of Maslow’s Hierarchy of Needs and Alderfer’s Existence-Relatedness-Growth (ERG) theory.
Figure 2: The illustration of the proposed innate-values-driven model.
However, natural agents, like humans, often make decisions based on a blend of biological, social, and cognitive motivations, as elucidated by combined motivations’ model like Maslow’s Hierarchy of Needs [6] and Alderfer’s Existence-Relatedness-Growth (ERG) theory [7]. Fig. 2 illustrates the five human agents with various personalities presenting different amounts of innate values and preferences based on five levels of Maslow’s Hierarchy of Needs and Alderfer’s ERG theory. On the other hand, the AI agent can be regarded as a self-organizing system that also presents various needs and motivations in its evolution through decision-making and learning to adapt to different scenarios and satisfy their needs [8].
@@
Many researchers regard motivated behavior as behavior that involves the assessment of the consequences of behavior through learned expectations, which makes motivation theories tend to be intimately linked to theories of learning and decision-making [9]. In particular, intrinsic motivation leads organisms to engage in exploration, play, strategies, and skills driven by expected rewards. The computational theory of reinforcement learning (RL) addresses how predictive values can be learned and used to direct behavior, making RL naturally relevant to studying motivation. For example, development RL is concerned with using deep RL algorithms to tackle a developmental problem – the intrinsically motivated acquisition of open-ended repertoires of skills [10].
@@
In artificial intelligence, researchers propose various abstract computational structures to form the fundamental units of cognition and motivations, such as states, goals, actions, and strategies. For intrinsic motivation modeling, the approaches can be generally classified into three categories: prediction-based [11, 12], novelty-based [13, 8], and competence-based [14, 15]. Furthermore, the concept of intrinsic motivation was introduced in machine learning and robotics to develop artificial systems learning diverse skills autonomously [16, 17, 18]. The idea is that intelligent machines and robots could autonomously acquire skills and knowledge under the guidance of intrinsic motivations and later exploit such knowledge and skills to accomplish tasks more efficiently and faster than if they had to acquire them from scratch [9].
@@
In other words, by investigating intrinsically motivated learning systems, we would clearly improve the utility and autonomy of intelligent artificial systems in dynamic, complex, and dangerous environments [19, 20]. Specifically, compared with the traditional RL model, intrinsically motivated RL refines it by dividing the environment into an external environment and an internal environment [21], which clearly generates all reward signals within the organism [^1] Here, the organism represents all the components of the internal environment in the AI agent. [9]. Although the extrinsic reward signals are triggered by the objects and events of the external environment, and activities of the internal environment cause the intrinsic reward signals, it is hard to determine the complexity and variability of the intrinsic rewards (innate values) generating mechanism. Specifically, traditional RL model is learning from interaction based on the environment’s feedback rewards. However, in real world, the rewards are generated by agents’ innate value systems, which differ vastly from individuals based on their needs and requirements. Moreover, the AI agent can be regarded as a self-organizing system that also presents various needs and motivations in its evolution through decision-making and learning to adapt to different scenarios and satisfy those needs. The traditional RL can not reasonably explain its innate values and motivations nor provide a long-term model to support the AI agent’s lifelong development.
@@
To address those gaps, we introduce the innate-values-driven reinforcement learning (IVRL) model, which integrates combined motivations’ models and expected utility theory to describe the complex behaviors in AI agents’ adaptation and evolution. We formalize the innate values and derive the IVRL model, then propose two corresponding algorithms: IV-DQN and IV-A2C. Furthermore, we compare them with benchmark RL algorithms such as DQN [22], DDQN [23], A2C [24], and PPO [25] in the Role-Playing Game (RPG) RL test platform VIZDoom [26, 27]. The results demonstrate that the proposed IVRL model can achieve convergence and adapt efficiently to complex and challenging tasks.
II Approach Overview
@@
We assume that all the AI agents (like robots) interact in the same working scenario, and their external environment includes all the other group members and mission setting. In contrast, the internal environment consists of individual perception components including various sensors (such as Lidar and camera), the critic module involving intrinsic motivation analysis and innate values generation, the RL brain making the decision based on the feedback of rewards and description of the current state (including internal and external) from the critic module, and actuators relating to all the manipulators and operators executing the RL brain’s decisions as action sequence and strategies. Fig. 2 illustrates the proposed innate-values-driven model.
Figure 3: The illustration of the IVRL model based on Expected Utility Theory.
@@
Compared with the traditional RL model, our model generates the input state and rewards from the critic module instead of directly from the environment, which means that the AI agent receives various utilities from the environment through executing an action or strategy in the IVRL model. Moreover, the individual needs to calculate innate values (expected utility) through its needs weights and current utilities and then select suitable actions or strategies to optimize or maximize its accumulated expected utility (Fig. 3). Specifically, we formalize the IVRL of an AI agent with an external environment using a Markov decision process (MDP) [28]. The MDP is defined by the tuple $\langle\mathcal{S,A,R,T,}\gamma\rangle$ where $\mathcal{S}$ represents the finite sets of internal state $S_{i}$ [^2] The internal state $S_{i}$ describes an agent’s innate value distribution and presents the dominant intrinsic motivation based on the external state $S_{e}$. and external states $S_{e}$. $\mathcal{A}$ represents a finite set of actions. The transition function $\mathcal{T}$: $\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow$ [0, 1] determines the probability of a transition from any state $s\in\mathcal{S}$ to any state $s^{\prime}\in\mathcal{S}$ given any possible action $a\in\mathcal{A}$. Assuming the critic function is $\mathcal{C}$, which describes the individual innate value model. The reward function $\mathcal{R}=\mathcal{C}(S_{e}):\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ defines the immediate and possibly stochastic innate reward $\mathcal{C}(S_{e})$ that an agent would receive given that the agent executes action $a$ which in state $s$ and it is transitioned to state $s^{\prime},\gamma\in[0,1)$ the discount factor that balances the trade-off between innate immediate and future rewards.
II-A The Expected Utility and Source of Randomness
In the IVRL model, the reward is regarded as the expected utility [29, 30] generated by the agent’s utility values $u(x_{k})$ and the corresponding probability $p_{k}$ ((1)). It is equal to the sum of each category’s needs weight $n_{k}$ times its current utility $u_{k}$ in the IVRL model ((4)).
$R_{t}=\sum_{i=1}^{k}u_{k}\times n_{k}=\mathbb{E}\left[U(p)\right]=\sum_{i=1}^{k}u(x_{k})p_{k}$ (1)
Furthermore, in the IVRL model, the randomness comes from three sources. The randomness in action is from the policy function: $A\sim\pi(\cdot|s)$; the needs weight function: $W\sim\omega(\cdot|s)$ makes the randomness of innate values; the state-transition function: $S^{\prime}\sim p(\cdot|s,a)$ causes the randomness in state.
@@ II-A The Expected Utility and Source of Randomness
Supposing at current state $s_{t}$ an agent has a needs weight matrix $N_{t}$ ((2)) in a mission, which presents its innate value weights for different levels of needs. Correspondingly, it has a utility matrix $U_{t}$ ((3)) for specific needs resulting from action $a_{t}$. Then, we can calculate its reward $R_{t}$ for $a_{t}$ through (4) at the state $s_{t}$.
In the process, the agent will first generate the needs weight and action based on the current state, then, according to the feedback utilities and the needs weights, calculate the current reward (expected utility) and iterate the process until the end of an episode. Fig. 4 illustrates the trajectory of state $S$, needs weight $W$, action $A$, and reward $R$, and Fig. 3 presents the corresponding IVRL model.
$N_{t}=\left[\begin{matrix}n_{11}&n_{12}&\cdots&n_{1m}\\ n_{21}&n_{22}&\cdots&n_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ n_{n1}&n_{n2}&\cdots&n_{nm}\\ \end{matrix}\right]$ (2)
$U_{t}=\left[\begin{matrix}u_{11}&u_{12}&\cdots&u_{1m}\\ u_{21}&u_{22}&\cdots&u_{2m}\\ \vdots&\vdots&\ddots&\vdots\\ u_{n1}&u_{n2}&\cdots&u_{nm}\\ \end{matrix}\right]$ (3)
@@
II-B Randomness in Discounted Returns
According to the above discussion, we define the discounted return $G$ at $t$ time as cumulative discounted rewards in the IVRL model ((5)) and $\gamma$ is the discount factor.
@@ II-B Randomness in Discounted Returns
II-C Action-Innate-Value Function
Based on the discounted return (5) and its random factors – (6), (7), and (8), we can define the Action-Innate-Value function as the expectation of the discounted return $G$ at $t$ time ((9)).
@@ II-C Action-Innate-Value Function
II-D State-Innate-Value Function
Furthermore, we can define the State-Innate-Value function as (10), which calculates the expectation of $Q_{\pi,\omega}(s_{t},w_{t},a_{t})$ for action $A$ and reflects the situation in the state $s_{t}$ with the innate value judgment $w_{t}$.
@@ II-D State-Innate-Value Function
II-E Approximate the Action-Innate-Value Function
The agent’s goal is to interact with the environment by selecting actions to maximize future rewards based on its innate value judgment. We make the standard assumption that a factor of $\gamma$ per time-step discounts future rewards and define the future discounted return at time $t$ as (5). Moreover, we can define the optimal action-value function $Q^{*}(s,a,w)$ as the maximum expected return achievable by following any strategy after seeing some sequence $s$, making corresponding innate value judgment $w$, and then taking action $a$, where $\omega$ is a needs weight function describing sequences about innate value weights and $\pi$ is a policy mapping sequences to actions.
@@ II-E Approximate the Action-Innate-Value Function
Furthermore, the same as the DQN [22], we use a function approximator ((13)) to estimate the action-innate-value function.
@@
Figure 5: Illustration of the IV-DQN network generating Needs-Behavior distribution.
@@
Where (15) is the target for iteration $i$ and $\sigma(s,w,a)$ is a probability distribution over sequences $s$, needs weights $w$, and action $a$ that we refer to as the needs-behavior distribution. We can approximate the gradient as follows:
$\begin{split}\nabla_{\theta_{i}}L_{i}(\theta_{i})=&\mathbb{E}_{s,w,a\sim\sigma(\cdot);s^{\prime}\sim\epsilon}\bigg{[}\bigg{(}r+\gamma\max_{w^{\prime},a^{\prime}}Q(s^{\prime},w^{\prime},a^{\prime};\theta_{i-1})\\ &-Q(s,w,a;\theta_{i})\bigg{)}\nabla_{\theta_{i}}Q(s,w,a;\theta_{i})\bigg{]}\end{split}$ (16)
Instead of computing the full expectations in the (16), stochastic gradient descent is usually computationally expedient to optimize the loss function. Here, the weights $\theta$ are updated after every time step, and single samples from the needs-behavior distribution $\sigma$ and the emulator $\epsilon$ replace the expectations, respectively.